OpenStudy (anonymous):

simplify and state restrictions x2-9/3-x

OpenStudy (anonymous):



OpenStudy (anonymous):

im not sure if this is right

OpenStudy (anonymous):

\[\frac{x^2-9}{3-x}=\frac{(x+3)(x-3)}{3-x}\]

OpenStudy (anonymous):

yeah it is right, but i am sure you math teacher would probably prefer \[-x-3\] as an answer

OpenStudy (anonymous):

oh okay thank you so much!!

OpenStudy (anonymous):

@satellite73 could you help me with one more im stuck on please?